OpenStudy (anonymous):

Prove that 5n^2 + 3n + 4 is even, for all integers n.

OpenStudy (anonymous):

i know that 2n is considered even and 2n + 1 is odd.. i tried to break it down (4n^2 + n^2) + (2n + n) + 4...

OpenStudy (thomas5267):

Do you know mathematical induction?

OpenStudy (anonymous):

learning it, haha

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 5n^2 + 3n + 4 }\) \(\large\color{black}{ \displaystyle 5n(n + 3) + 4 }\) when 'n' is even you get an even 5n times an odd n+3 which is an even number. Then add 4 to this even number -this is still even. when 'n' is odd, then n+3 is even and 5n is odd, and even n+3 times odd 5n is an even number. Again, add 4 to this even number, and your output is still going to be even. this is what I would say.

OpenStudy (solomonzelman):

or, instead of analyzing you can break it down as you did, \(\large\color{black}{ \displaystyle 5n^2 + 3n + 4 }\) \(\large\color{black}{ \displaystyle 4n^2 + n^2+2n+n + 4 }\) \(\large\color{black}{ \displaystyle 4n^2 +2n+ n^2+n + 4 }\) \(\large\color{black}{ \displaystyle 2n(2n +1)+ n(n+1) + 4 }\) and here, again regardless if you choose an odd or even "n' (as long as it is integer) you get an even output

OpenStudy (anonymous):

its beautiful

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 2n(2n+1)~~~~~~~~+~n(n+1)~~~+4 }\) \(\large\color{black}{ \displaystyle ({\rm even } \times {\rm odd})~~~~+{\rm even } \times {\rm odd}~~~+{\rm even } }\) OR, when n is an odd number \(\large\color{black}{ \displaystyle ({\rm even } \times {\rm odd})~~~~+{\rm odd } \times {\rm even}~~~+{\rm even } }\)

OpenStudy (solomonzelman):

(don't want to do anything complicated on this easy question)

OpenStudy (solomonzelman):

but, you get the point.....

OpenStudy (anonymous):

thanks, yeah i just have a hard time developing a "proof-like" answer still, need more practice! Thank you

OpenStudy (solomonzelman):

sure... yw !